Kai receives a 5% commission on the first $6000 of sales and 7% for sales over $6000.
If the commission is $795, how much was sold?
my work: 5% of $6000= $300
this was as far as i could get.
thank you for your help!

Answer by richwmiller(17219)

6000+y=x
x is total sales
.05*6000=300 commission on first 6000
795-300=495 commission on sales over 6000
.07*y=495
y=495/.07
y=7071.43 sales over 6000
x=13071.43 total sales